# thanks: Full author list given at the end of the paper

### v2

@@ Abstract
The study of the charmed baryons is crucial for investigating the strong and weak interactions in the Standard Model and for gaining insights into the internal structure of baryons. In an $e^{+}e^{-}$ experiment the lightest charmed baryon, $\Lambda_{c}^{+}$, can be produced in pairs through the single photon annihilation process. This process can be described by two complex electromagnetic form factors. The presence of a non-zero relative phase between these form factors gives rise to a transverse polarization of the charmed baryon and provides additional constraints on the dynamic parameters in the decays. In this article, we present the first observation of the transverse polarization of $\Lambda_{c}^{+}$ in the reaction $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, based on $6.4~\text{fb}^{-1}$ of $e^{+}e^{-}$ annihilation data collected at center-of-mass energies between $4600$ MeV and $4951$ MeV with the BESIII detector. The decay asymmetry parameters and strong phase shift in the decays $\Lambda_{c}^{+}\to pK_{S}^{0}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, $\Sigma^{+}\pi^{0}$ are also simultaneously extracted from the joint angular distributions. These results are vital for understanding CPC\!P violation and its role in the matter-antimatter asymmetry of the Universe.
@@ I Introduction
Although the most abundant constituent of matter is baryons, our understanding of them is less knowledgeable than that of mesons. The addition of one constituent quark in baryons compared to the quark-antiquark mesons introduces numerous challenges in theoretical calculations, making high-order and diquark correlation contributions difficult to predict accurately. One of the effective approaches to study the structure of baryons is to measure their electromagnetic form factors (EMFFs) [1]. In the $e^{+}e^{-}$ annihilation process, the time-like form factors are accessible, providing insights into the electromagnetic structure of the baryons [2]. For a spin-1/2 baryon ($B$), the process of $e^{+}e^{-}\to B\bar{B}$ is dominated by the single photon exchange. The cross-section can be described by the complex electric form factor $G_{E}$ and magnetic form factor $G_{M}$, which are functions of the four-momentum squared $q^{2}$ of the photon. The angular distribution of the produced baryon is characterized by two parameters: the angular parameter $\alpha_{0}$ and the phase difference between the EMFFs $\Delta\Phi$, where $\Delta\Phi$ is defined as $\text{arg}(G_{M})-\text{arg}(G_{E})$ [3]. The transverse polarization can be parameterized as
$P_{y}(\cos\theta_{0})=\frac{3}{2(3+\alpha_{0})}\sqrt{1-\alpha^{2}_{0}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi,$ (1)
@@
where the factor $\frac{3}{2(3+\alpha_{0})}$ arises from normalization in spin density matrix of $\Lambda_{c}^{+}$. A non-vanishing $\Delta\Phi$ gives rise to the transverse polarization effect of the baryon, which depends on the momentum direction of the baryon ($\theta_{0}$) in the rest frame of $e^{+}e^{-}$. This polarization effect has been observed in the pair production of hyperons $\Lambda$ [4, 5, 6], $\Sigma^{0,+}$ [7, 8, 9] and $\Xi^{0,-}$ [10, 11, 12] by the BESIII Collaboration, which has led to the highest precision measurements of the decay parameters [13] of the hyperons. However, this effect has never been observed in charmed baryons.
The $\Lambda_{c}^{+}$ is the lightest charmed baryon, where a light up quark in the proton is replaced by a heavy charm quark. The study of the $\Lambda_{c}^{+}$ decay is of significant importance in investigating the fundamental theory of the strong and weak interactions. Similar to the pair production of strange hyperons, it is expected that the $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ production can also acquire polarization with respect to the normal of the production plane.
Figure 1: Generation and transmission of polarization in a weak decay $\Lambda_{c}^{+}\to BP$ [14].
@@
In the weak decays of $\Lambda_{c}^{+}\to BP$, where $B$ and $P$ denote a spin-1/2 baryon and a pseudoscalar meson, respectively, the polarization of $\Lambda_{c}^{+}$, $\bm{\mathcal{P}_{\Lambda_{c}^{+}}}$, can be transferred to the subsequent decays and manifested in the angular distributions of the daughter baryons, as shown in Fig. 1. The polarization of the daughter baryon, denoted by $\bm{\mathcal{P}_{B}}$, is expressed as
@@
$\small\bm{\mathcal{P}_{B}}=\frac{(\alpha+\bm{\mathcal{P}_{\Lambda_{c}^{+}}}\cdot\bm{\hat{n}})\bm{\hat{n}}+\beta(\bm{\mathcal{P}_{\Lambda_{c}^{+}}}\times\bm{\hat{n}})+\gamma\bm{\hat{n}}\times(\bm{\mathcal{P}_{\Lambda_{c}^{+}}}\times\bm{\hat{n}})}{1+\alpha\bm{\mathcal{P}_{B}}\cdot\bm{\hat{n}}}$, (2)
[…]
@@
Various phenomenological models taking into account non-perturbative effects and non-factorizable contributions in quantum chromodynamics (QCD), have been developed to predict the decay asymmetries of the charmed baryons. The development of these models can be roughly divided into three stages. The first stage was in the last century, after CLEO Collaboration first measured $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ in 1990 [16]. Models such as covariant confined quark model (CCQM) [17, 18], the pole model [19, 20, 21, 22, 23], and current algebra (CA) [24, 25] were applied to predict decay asymmetry parameters. The decay asymmetry parameters $\alpha_{pK^{0}_{S}}$, $\alpha_{\Lambda\pi^{+}}$, $\alpha_{\Sigma^{0}\pi^{+}}$, and $\alpha_{\Sigma^{+}\pi^{0}}$ predicted by different models varied significantly, particularly with regard to uncertainty on the sign. The beginning of the second stage is marked by the BESIII Collaboration’s measurement of the decay asymmetry in $\Lambda_{c}^{+}\to pK^{0}_{S}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ decays together with the first determination of $\alpha_{pK^{0}_{S}}$ and $\alpha_{\Sigma^{0}\pi^{+}}$ in 2019 [15]. The sign of $\alpha_{pK^{0}_{S}}$ is positive with a large uncertainty, while all predictions are negative. Another point is that the sign of $\alpha_{\Sigma^{0}\pi^{+}}$ has been confirmed. The Belle Collaboration results for $\alpha_{\Sigma^{0}\pi^{+}}$ and $\alpha_{\Sigma^{+}\pi^{0}}$ [26, 27] are in a good agreement with the flavor SU(3) symmetry. Subsequently, the BESIII Collaboration first reported the $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ decay asymmetry parameters in 2024 [28], which is a pure $W$-exchange process. The non-perturbative nature and the problem of non-factorization in this decay channel make predictions very difficult. A persistent tension exists between the branching fraction and the decay asymmetry parameters, as theoretical predictions and experimental measurements do not align perfectly. To resolve this tension, the study of the $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ decay introduces a valuable perspective as the strong phase shift, which describes the relative phase difference induced by strong interactions in the decay, has a crucial role and may significantly influence the decay asymmetry parameters. Following this, some theoretical works [29, 30, 31] have provided detailed discussions on this measurement and also introduced a strong phase shift, which significantly improves the description of the data. The most notable change due to the strong phase shift is seen for $\alpha_{pK^{0}_{S}}$, where this parameter changes from a negative value to one consistent with zero. This modification leads to a substantial change in the related calculations, making the precise measurement of $\alpha_{pK^{0}_{S}}$ crucial. In 2024, the LHCb Collaboration measured $\alpha_{pK^{0}_{S}}$ and $\alpha_{\Lambda\pi^{+}}$ with high accuracy using the $\Lambda_{b}^{0}\to\Lambda_{c}^{+}\pi^{-}$ decay, which provides a large longitudinal polarization to $\Lambda_{c}^{+}$ in the decay process  [32]. These results agree with the latest theoretical predictions  [33]. It is clear that the value of $\alpha_{pK^{0}_{S}}$ should be negative rather than close to zero. The BESIII Collaboration can also independently measure $\alpha_{pK^{0}_{S}}$ using a large sample to confirm the sign of $\alpha_{pK^{0}_{S}}$. For $\alpha_{\Lambda\pi^{+}}$, the Belle Collaboration reported $\alpha_{\Lambda\pi^{+}}$ [26] with the same accuracy as the result from the LHCb Collaboration, but did not report $\beta_{\Lambda\pi^{+}}$ and $\gamma_{\Lambda\pi^{+}}$. All theoretical predictions and measurements about $\alpha_{pK^{0}_{S}}$, $\alpha_{\Lambda\pi^{+}}$, $\alpha_{\Sigma^{0}\pi^{+}}$, and $\alpha_{\Sigma^{+}\pi^{0}}$ are summarized in Table 1.
@@
Predictions or experiments $\alpha_{pK^{0}_{S}}$ $\alpha_{\Lambda\pi^{+}}$ $\alpha_{\Sigma^{0}\pi^{+}}$ $\alpha_{\Sigma^{+}\pi^{0}}$
Körner(1992), CCQM [17] 0.10-0.10 0.70-0.70 $0.70$ $0.71$
Xu(1992), Pole [19] $0.51$ 0.67-0.67 $0.92$ $0.92$
Cheng, Tseng(1992), Pole [20] 0.49-0.49 0.96-0.96 $0.83$ $0.83$
Cheng, Tseng(1993), Pole [21] 0.49-0.49 0.95-0.95 $0.78$ $0.78$
Źencaykowski(1994), Pole [22] 0.66-0.66 0.99-0.99 $0.39$ $0.39$
Źencaykowski(1994), Pole [23] 0.90-0.90 0.86-0.86 0.76-0.76 0.76-0.76
Alakabha Datta(1995), CA [24] 0.91-0.91 0.94-0.94 0.47-0.47 0.47-0.47
Ivanov(1998), CCQM [18] 0.97-0.97 0.95-0.95 $0.43$ $0.43$
Sharma(1999), CA [25] 0.99-0.99 0.99-0.99 0.31-0.31 0.31-0.31
Geng(2019), SU(3) [34] $0.89^{+0.26}_{-0.11}$ $0.87\pm 0.10$ $0.35\pm 0.27$ $0.35\pm 0.27$
Zou(2020), CA [35] 0.75-0.75 0.93-0.93 0.76-0.76 0.76-0.76
Zhong (2022), SU(3)a [36] $0.51\pm 0.21$ $0.75\pm 0.01$ $0.47\pm 0.03$ $0.47\pm 0.03$
Zhong (2022), SU(3)b [36] $0.29\pm 0.24$ $0.75\pm 0.01$ $0.47\pm 0.03$ $0.47\pm 0.03$
Liu(2023), Pole [37] $0.81\pm 0.05$ $0.75\pm 0.01$ $0.47\pm 0.01$ $0.45\pm 0.04$
Liu(2023), LP [37] $0.68\pm 0.01$ $0.75\pm 0.01$ $0.47\pm 0.01$ $0.45\pm 0.04$
Geng(2023), SU(3) [29] $0.40\pm 0.49$ $0.75\pm 0.01$ $0.47\pm 0.02$ $0.47\pm 0.02$
Zhong(2024), TDA [31] $0.01\pm 0.24$ $0.76\pm 0.01$ $0.48\pm 0.02$ $0.48\pm 0.02$
Zhong(2024), IRA [31] $0.03\pm 0.24$ $0.76\pm 0.01$ $0.48\pm 0.02$ $0.48\pm 0.02$
Zhong(2024), TDA [33] $0.74\pm 0.03$ $0.76\pm 0.01$ $0.47\pm 0.01$ $0.47\pm 0.01$
Zhong(2024), IRA [33] $0.74\pm 0.03$ $0.76\pm 0.01$ $0.47\pm 0.01$ $0.47\pm 0.01$
CLEO(1990) [16] $1.0^{+0.4}_{-0.1}$
ARGUS(1992) [38] $0.96\pm 0.42$
CLEO(1995) [39] $0.94^{+0.21+0.12}_{-0.06-0.06}$ $0.45\pm 0.31\pm 0.06$
FOCUS(2006) [40] $0.78\pm 0.16\pm 0.19$
BESIII(2019) [15] $0.18\pm 0.43\pm 0.14$ $0.80\pm 0.11\pm 0.02$ $0.73\pm 0.17\pm 0.07$ $0.57\pm 0.10\pm 0.07$
BELLE(2022) [26, 27] $0.755\pm 0.005\pm 0.003$ $0.463\pm 0.016\pm 0.008$ $0.48\pm 0.02\pm 0.02$
LHCb(2024) [32] $0.754\pm 0.008\pm 0.006$ $0.785\pm 0.006\pm 0.003$
PDG Fit(2025) [41] $0.754\pm 0.010$ $0.768\pm 0.015$ $0.466\pm 0.018$ $0.484\pm 0.027$
@@
The asymmetry in charge conjugation and parity symmetry (CPC\!P) lies at the heart of some of the most fundamental problems in physics. The presence of CPC\!P asymmetry has been established in meson decays [42, 43, 44]. Recently, the LHCb collaboration reported the first observation of direct CPC\!P asymmetry in a baryonic decay [45], $\Lambda_{b}^{0}\to pK^{-}\pi^{+}\pi^{-}$, marking a significant milestone in the study of CPC\!P violation. Meanwhile, ongoing efforts are exploring CPC\!P asymmetry in other baryonic processes [46, 47]. Despite these discoveries, the origin of CPC\!P asymmetry and a precise determination of its magnitude in specific processes remain open questions. Its size strictly depends on the strong and weak phase shift. In Cabbibo-favored (CF) decays of the $\Lambda_{c}^{+}$ baryon, there is no CPC\!P asymmetry, as the weak phase difference is expected to be zero. By relaxing the constraints of CPC\!P asymmetry , we can independently extract both the strong and weak phase shifts. The strong phase shift is an important input for theoretical calculations, helping to refine models of hadronic dynamics and improve our understanding of the underlying strong interaction. On the other hand, the weak phase provides a direct test of Standard Model (SM) predictions, offering valuable insights into potential CPC\!P asymmetry and verifying the consistency of weak interactions in charm decays.

### v3

@@ Abstract
The study of the charmed baryons is crucial for investigating the strong and weak interactions in the Standard Model and for gaining insights into the internal structure of baryons. In an $e^{+}e^{-}$ experiment the lightest charmed baryon, $\Lambda_{c}^{+}$, can be produced in pairs through the single photon annihilation process. This process can be described by two complex electromagnetic form factors. The presence of a non-zero relative phase between these form factors gives rise to a transverse polarization of the charmed baryon and provides additional constraints on the dynamic parameters in the decays. In this article, we present the first observation of the transverse polarization of $\Lambda_{c}^{+}$ in the reaction $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$, based on $6.4~\text{fb}^{-1}$ of $e^{+}e^{-}$ annihilation data collected at center-of-mass energies between $4600$ MeV and $4951$ MeV with the BESIII detector. The decay asymmetry parameters in the decays $\Lambda_{c}^{+}\to pK_{S}^{0}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ are simultaneously extracted from the joint angular distributions. From these parameters, both the weak and strong phase shifts are extracted, and several CPC\!P observables are tested. The obtained results show that the transverse polarization of the produced baryon significantly enhances the capability of searching for CPC\!P-violating effects, opening a novel way for exploring CPC\!P violation in the charmed baryon sector.
@@ I Introduction
Although baryons populate with a significant abundance in the physical world surrounding us, our understanding of them is less knowledgeable than that of mesons. The addition of one constituent quark in baryons compared to the quark-antiquark mesons introduces numerous challenges in theoretical calculations, making high-order and diquark correlation contributions difficult to predict accurately. One of the effective approaches to study the structure of baryons is to measure their electromagnetic form factors (EMFFs) [Schonning:2023mge, Huang:2021xte, Pacetti:2014jai]. In the $e^{+}e^{-}$ annihilation process, the time-like form factors are accessible, providing insights into the electromagnetic structure of the baryons [Huang:2021xte]. For a spin-1/2 baryon ($B$), the process of $e^{+}e^{-}\to B\bar{B}$ is dominated by the single photon exchange. The cross-section can be described by the complex electric form factor $G_{E}$ and magnetic form factor $G_{M}$, which are functions of the four-momentum squared $q^{2}$ of the photon. The angular distribution of the produced baryon is characterized by two parameters: the angular parameter $\alpha_{0}$ and the phase difference between the EMFFs $\Delta\Phi$, where $\Delta\Phi$ is defined as $\text{arg}(G_{M})-\text{arg}(G_{E})$ [Chen:2019hqi]. The transverse polarization can be parameterized as
$\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}(\cos\theta_{0})=\frac{3}{2(3+\alpha_{0})}\sqrt{1-\alpha^{2}_{0}}\sin\theta_{0}\cos\theta_{0}\sin\Delta\Phi,$ (1)
@@
where the factor $\frac{3}{2(3+\alpha_{0})}$ arises from normalization in spin density matrix of $\Lambda_{c}^{+}$. The $\alpha_{0}$ can be written as $\alpha_{0}=(|G_{E}|^{2}(1-v^{2})-|G_{M}|^{2})/(|G_{E}|^{2}(1-v^{2})+|G_{M}|^{2})$, where the $v$ is the velocity of the $\Lambda_{c}^{+}$ in center-of-mass (CM) system, normalized to the speed of light. A non-vanishing $\Delta\Phi$ gives rise to the transverse polarization effect of the baryon, which depends on the momentum direction of the baryon ($\theta_{0}$) in the rest frame of $e^{+}e^{-}$, as shown in Fig. 1.
Figure 1: Definition of the coordinate system and kinematic variables for the process $e^{+}e^{-}\to\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$. The production angle $\theta_{0}$ is measured between the positron beam and the $\Lambda_{c}^{+}$ momentum in the center-of-mass frame, and the polarization $\mathcal{P}_{\Lambda_{c}^{+}}^{y}$ is oriented perpendicular to the production plane.
This polarization effect has been observed in the pair production of hyperons $\Lambda$ [BESIII:2018cnd, BESIII:2022qax, BESIII:2023euh], $\Sigma^{0,+}$ [BESIII:2023cvk, BESIII:2023ynq, BESIII:2024nif] and $\Xi^{0,-}$ [BESIII:2021ypr, BESIII:2022lsz, BESIII:2023drj] by the BESIII Collaboration, which has led to the highest precision measurements of the decay parameters [Lee:1957qs] of the hyperons. Similar to the pair production of hyperons, it is expected that the $\Lambda_{c}^{+}$ via $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ production can acquire polarization with respect to the normal of the production plane. However, this effect has never been established in charmed baryons.
Investigations of charmed baryon decay dynamics are essential for exploring the weak and strong interactions within the standard model (SM) of particle physics. The lightest charmed baryon, $\Lambda_{c}^{+}$, was discovered in 1979 [Abrams:1979iu]. It consists of one heavy charm quark and two light $u$ and $d$ quarks, being bound by the strong interaction. The charm quark mass lies in the intermediate region between perturbative and non-perturbative quantum chromodynamics (QCD), making the study of charmed baryons crucial for understanding non-perturbative effects in QCD.
@@
In the weak decays of $\Lambda_{c}^{+}\to BP$, where $B$ and $P$ denote a spin-1/2 baryon and a pseudoscalar meson, respectively, the polarization of $\Lambda_{c}^{+}$, $\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}$, can be transferred to the subsequent decays and manifested in the angular distributions of the daughter baryons, as shown in Fig. 2.
Figure 2: Generation and transfer of polarization in the weak decay $\Lambda_{c}^{+}\to BP$ [Wang:2024wrm]. The polarization of the daughter baryon $\bm{\mathcal{P}_{B}}$ depends on that of the parent $\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}$, the decay angle $\theta$, and the weak decay parameters $\alpha$, $\beta$, and $\gamma$.
The polarization of the daughter baryon, denoted by $\bm{\mathcal{P}_{B}}$, is expressed as
@@
$\small\bm{\mathcal{P}_{B}}=\frac{(\alpha+\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}\cdot\bm{\hat{n}})\bm{\hat{n}}+\beta(\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}\times\bm{\hat{n}})+\gamma\bm{\hat{n}}\times(\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}\times\bm{\hat{n}})}{1+\alpha\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y}}\cdot\bm{\hat{n}}}$, (2)
where $\alpha$, $\beta$, and $\gamma$ are the decay asymmetry parameters of $\Lambda_{c}^{+}\to BP$ and $\bm{\hat{n}}$ is the unit vector along the baryon flight direction in the rest frame of $\Lambda_{c}^{+}$. Previous study of the transverse polarization of the $\Lambda_{c}^{+}$ using data collected at a CM energy of $4600$ MeV at BESIII, only had a statistical significance of 2.1 standard deviations [BESIII:2019odb].
@@
Various phenomenological models taking into account non-perturbative effects and non-factorizable contributions in QCD, have been developed to calculate the decay asymmetries of the charmed baryons. The development of these models can be roughly divided into three stages. The first stage was in the last century, after the CLEO Collaboration first measured $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ in 1990 [CLEO:1990unw]. Models such as covariant confined quark model (CCQM) [Korner:1992wi, Ivanov:1997ra], the pole model [Xu:1992vc, Cheng:1991sn, Cheng:1993gf, Zenczykowski:1993jm, Zenczykowski:1993hw], and current algebra (CA) [Datta:1995mn, Sharma:1998rd] were applied to calculate decay asymmetry parameters. The decay asymmetry parameters $\alpha_{pK^{0}_{S}}$, $\alpha_{\Lambda\pi^{+}}$, $\alpha_{\Sigma^{0}\pi^{+}}$, and $\alpha_{\Sigma^{+}\pi^{0}}$ calculated by different models varied significantly, particularly with regard to uncertainty on the sign. The beginning of the second stage is marked by the BESIII Collaboration’s measurement of the decay asymmetry in $\Lambda_{c}^{+}\to pK^{0}_{S}$, $\Lambda\pi^{+}$, $\Sigma^{0}\pi^{+}$, and $\Sigma^{+}\pi^{0}$ decays together with the first determination of $\alpha_{pK^{0}_{S}}$ and $\alpha_{\Sigma^{0}\pi^{+}}$ in 2019 [BESIII:2019odb]. The central value of $\alpha_{pK^{0}_{S}}$ is positive with a large uncertainty, and most of the predictions were negative. Thus, the sign of $\alpha_{pK^{0}_{S}}$ remained a puzzle at that time. Another point is that the sign of $\alpha_{\Sigma^{0}\pi^{+}}$ has been confirmed. The Belle Collaboration results for $\alpha_{\Sigma^{0}\pi^{+}}$ and $\alpha_{\Sigma^{+}\pi^{0}}$ [Belle:2022uod, Belle:2022bsi] are in a good agreement with the flavor SU(3) symmetry. Subsequently, the BESIII Collaboration first reported the $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ decay asymmetry parameters in 2024 [BESIII:2023wrw], which is a pure $W$-exchange process. The non-perturbative nature and the problem of non-factorization in this decay channel make calculations very difficult. A persistent tension exists between the branching fraction and the decay asymmetry parameters, as theoretical calculations and experimental measurements do not align perfectly. To resolve this tension, the study of the $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ decay introduces a valuable perspective as the strong phase shift, which describes the relative phase difference induced by strong interactions in the decay, has a crucial role and may significantly influence the decay asymmetry parameters. Following this, some theoretical works [Geng:2023pkr, Zhong:2024zme, Zhong:2024qqs] have provided detailed discussions on this measurement and also introduced a strong phase shift, which significantly improves the description of the data. The most notable change due to the strong phase shift is seen for $\alpha_{pK^{0}_{S}}$, whose value becomes close to 0 [Zhong:2024qqs]. This modification leads to a substantial change in the related calculations, making the precise measurement of $\alpha_{pK^{0}_{S}}$ crucial. In 2024, the LHCb Collaboration measured $\alpha_{pK^{0}_{S}}$ and $\alpha_{\Lambda\pi^{+}}$ with high accuracy using the $\Lambda_{b}^{0}\to\Lambda_{c}^{+}\pi^{-}$ decay, which provides a large longitudinal polarization to $\Lambda_{c}^{+}$ in the decay process  [LHCb:2024tnq]. These results agree with the latest theoretical calculations  [Cheng:2024lsn]. It is clear that the value of $\alpha_{pK^{0}_{S}}$ should be negative rather than close to zero. The BESIII Collaboration can also independently measure $\alpha_{pK^{0}_{S}}$ using a large sample to confirm the sign of $\alpha_{pK^{0}_{S}}$. For $\alpha_{\Lambda\pi^{+}}$, the Belle Collaboration reported $\alpha_{\Lambda\pi^{+}}$ [Belle:2022uod] with the same accuracy as the result from the LHCb Collaboration, but did not report $\beta_{\Lambda\pi^{+}}$ and $\gamma_{\Lambda\pi^{+}}$. All theoretical calculations and measurements about $\alpha_{pK^{0}_{S}}$, $\alpha_{\Lambda\pi^{+}}$, $\alpha_{\Sigma^{0}\pi^{+}}$, and $\alpha_{\Sigma^{+}\pi^{0}}$ are summarized in Table 1.
@@
Calculations or experiments $\alpha_{pK^{0}_{S}}$ $\alpha_{\Lambda\pi^{+}}$ $\alpha_{\Sigma^{0}\pi^{+}}$ $\alpha_{\Sigma^{+}\pi^{0}}$
Körner(1992), CCQM [Korner:1992wi] 0.10-0.10 0.70-0.70 $0.70$ $0.71$
Xu(1992), Pole [Xu:1992vc] $0.51$ 0.67-0.67 $0.92$ $0.92$
Cheng, Tseng(1992), Pole [Cheng:1991sn] 0.49-0.49 0.96-0.96 $0.83$ $0.83$
Cheng, Tseng(1993), Pole [Cheng:1993gf] 0.49-0.49 0.95-0.95 $0.78$ $0.78$
Źencaykowski(1994), Pole [Zenczykowski:1993jm] 0.66-0.66 0.99-0.99 $0.39$ $0.39$
Źencaykowski(1994), Pole [Zenczykowski:1993hw] 0.90-0.90 0.86-0.86 0.76-0.76 0.76-0.76
Alakabha Datta(1995), CA [Datta:1995mn] 0.91-0.91 0.94-0.94 0.47-0.47 0.47-0.47
Ivanov(1998), CCQM [Ivanov:1997ra] 0.97-0.97 0.95-0.95 $0.43$ $0.43$
Sharma(1999), CA [Sharma:1998rd] 0.99-0.99 0.99-0.99 0.31-0.31 0.31-0.31
Geng(2019), SU(3) [Geng:2019xbo] $0.89^{+0.26}_{-0.11}$ $0.87\pm 0.10$ $0.35\pm 0.27$ $0.35\pm 0.27$
Zou(2020), CA [Zou:2019kzq] 0.75-0.75 0.93-0.93 0.76-0.76 0.76-0.76
Zhong (2022), SU(3)a [Zhong:2022exp] $0.51\pm 0.21$ $0.75\pm 0.01$ $0.47\pm 0.03$ $0.47\pm 0.03$
Zhong (2022), SU(3)b [Zhong:2022exp] $0.29\pm 0.24$ $0.75\pm 0.01$ $0.47\pm 0.03$ $0.47\pm 0.03$
Liu(2023), Pole [Liu:2023dvg] $0.81\pm 0.05$ $0.75\pm 0.01$ $0.47\pm 0.01$ $0.45\pm 0.04$
Liu(2023), LP [Liu:2023dvg] $0.68\pm 0.01$ $0.75\pm 0.01$ $0.47\pm 0.01$ $0.45\pm 0.04$
Geng(2023), SU(3) [Geng:2023pkr] $0.40\pm 0.49$ $0.75\pm 0.01$ $0.47\pm 0.02$ $0.47\pm 0.02$
Zhong(2024), TDA [Zhong:2024qqs] $0.01\pm 0.24$ $0.76\pm 0.01$ $0.48\pm 0.02$ $0.48\pm 0.02$
Zhong(2024), IRA [Zhong:2024qqs] $0.03\pm 0.24$ $0.76\pm 0.01$ $0.48\pm 0.02$ $0.48\pm 0.02$
Zhong(2024), TDA [Cheng:2024lsn] $0.74\pm 0.03$ $0.76\pm 0.01$ $0.47\pm 0.01$ $0.47\pm 0.01$
Zhong(2024), IRA [Cheng:2024lsn] $0.74\pm 0.03$ $0.76\pm 0.01$ $0.47\pm 0.01$ $0.47\pm 0.01$
CLEO(1990) [CLEO:1990unw] $1.0^{+0.4}_{-0.1}$
ARGUS(1992) [ARGUS:1991yzs] $0.96\pm 0.42$
CLEO(1995) [CLEO:1995qyd] $0.94^{+0.21+0.12}_{-0.06-0.06}$ $0.45\pm 0.31\pm 0.06$
FOCUS(2006) [FOCUS:2005vxq] $0.78\pm 0.16\pm 0.19$
BESIII(2019) [BESIII:2019odb] $0.18\pm 0.43\pm 0.14$ $0.80\pm 0.11\pm 0.02$ $0.73\pm 0.17\pm 0.07$ $0.57\pm 0.10\pm 0.07$
BELLE(2022) [Belle:2022uod, Belle:2022bsi] $0.755\pm 0.005\pm 0.003$ $0.463\pm 0.016\pm 0.008$ $0.48\pm 0.02\pm 0.02$
LHCb(2024) [LHCb:2024tnq] $0.754\pm 0.008\pm 0.006$ $0.785\pm 0.006\pm 0.003$
PDG Fit(2025) [ParticleDataGroup:2024cfk] $0.754\pm 0.010$ $0.768\pm 0.015$ $0.466\pm 0.018$ $0.484\pm 0.027$
@@
The asymmetry in charge conjugation and parity symmetry (CPC\!P) lies at the heart of some of the most fundamental problems in physics. The presence of CPC\!P asymmetry has been established in meson decays [KTeV:1999kad, LHCb:2019hro, BaBar:2004gyj]. In 2019, the LHCb Collaboration observed a difference in time-integrated CPC\!P asymmetries between the decays $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ [LHCb:2019hro], providing the first observation of direct CPC\!P violation in the charm sector. This established that all three quark sectors, strange, bottom, and charm, exhibit CPC\!P-violating effects, albeit at very different magnitudes. Recently, the LHCb Collaboration reported the first observation of direct CPC\!P asymmetry in a baryonic decay [LHCb:2025ray], $\Lambda_{b}^{0}\to pK^{-}\pi^{+}\pi^{-}$, marking a significant milestone in the study of CPC\!P violation. Meanwhile, ongoing efforts are exploring CPC\!P asymmetry in other baryonic processes [LHCb:2016yco, LHCb:2024yzj]. Despite these discoveries, the origin of CPC\!P asymmetry and a precise determination of its magnitude in specific processes remain open questions. In the SM, CPC\!P violation arises from the interplay between strong and weak phases: weak phases change sign under a CPC\!P transformation, while strong phases remain invariant. A measurement of the CPC\!P asymmetry thus enables extraction of these two phases, and if one is known, the other can be determined. The strong phase shift is an important input for theoretical calculations, helping to refine models of hadronic dynamics and improve our understanding of the underlying strong interaction. On the other hand, the weak phase provides a direct test of SM predictions, offering valuable insights into potential CPC\!P asymmetry and verifying the consistency of weak interactions in charm decays. In Cabibbo-favored (CF) decays of the $\Lambda_{c}^{+}$, there is no CPC\!P asymmetry in SM, as the weak phase difference is expected to be zero. Although observing a non-zero weak phase shift in these processes is currently unrealistic based on the current data sample, a well-measured strong phase shift would still enhance the potential observability of a weak phase shift.